OpenStudy (anonymous):

f(x)= 4e^(xcos(x)) deravative? please

hartnn (hartnn):

knpow about chain rule ?

OpenStudy (anonymous):

yeah but wouldn't ex be the same

hartnn (hartnn):

\(\Large (4e^{x \cos x})' = 4e^{x\cos x} \times (x\cos x )'\)

OpenStudy (anonymous):

what would the derivative of the (xcosx )be? please tell me!

OpenStudy (anonymous):

by chain rule

hartnn (hartnn):

know product rule ?

hartnn (hartnn):

\((fg)' = f'g+fg'\)

hartnn (hartnn):

here f= x g = cos x

OpenStudy (anonymous):

ohh right haha, thanku

OpenStudy (anonymous):

sorry I mean product rule

OpenStudy (anonymous):

\[der(x)=4e ^{xcosx}[x(sinx)+cosx]\]

OpenStudy (anonymous):

[cosx - xsinx]

hartnn (hartnn):

yes :)

OpenStudy (anonymous):

THankyou!! guys

hartnn (hartnn):

welcome girl ^_^

OpenStudy (anonymous):

most welcome